Zinc metal reacts with hydrochloric acid according to the following balanced equation:

Zn(s)+2HCl(aq)→ ZnCl2(aq)+H2(g)

When 0.105 g of Zn(s) is combined with enough HCl to make 50.4 mL of solution in a coffee-cup calorimeter, all of the zinc reacts, raising the temperature of the solution from 22.2 ∘C to 24.7 ∘C.

Find ΔHrxn for this reaction as written. (Use 1.00 g/mL for the density of the solution and 4.18 J/(g⋅∘C) as the specific heat capacity.)

Express your answer in kilojoules per mole to two significant figures

(show work please)
Answer #1

Solution-

The Mass of the solution can be find as = volume x density

= 50.4 x 1.0 = 50.4 g

So the heat released by reaction = heat absorbed by solution

= mass x specific heat x temperature change of solution

= 50.4 * 4.18 * (24.7 - 22.4)

= 484.54 J

Now the moles of Zn = mass/molar mass of Zn

= 0.105/65.41 = 0.0016 mol

Let's now find ΔH = -heat released by reaction/moles of Zn

= -484.54/0.0016

= -3.02 * 10^5 J/mol

= -302 kJ/mol

Here ΔH is negative because the reaction is exothermic and heat is released